Female 95 91 8574 7580 79 8784 9686 100 C. Siance the P t value s 005, there is no Are puse ates affected by gender? A Snce the Palue for gender in argera ane affected by gender 무 v1/2010 2 4 8 9 tab :// Q 11 W/ 11 E || R || Tİl Y. || U
Answer #1

Solution:-

SUMMARY Under 20 20 to 40 Over 40 Total Male Count Sum Average Variance 12 937 83 78.08333 70.25 20.6666751.33333 56.99242 31

a) Since p-value is greater than 0.05, there is no significant interaction.

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: There is no interaction between gender and age.

Alternative hypothesis: There is interaction between gender and age.

Formulate an analysis plan. For this analysis, the significance level is 0.05.

Analyze sample data.

F statistics is given by:-

F=\frac{MSB}{MSE}

F = 0.148

Fcritical = 3.56

The P-value = 0.86

Interpret results. Since the P-value (0.86) is greater than the significance level (0.05), we have to accept the null hypothesis.

Conclusion:-

Do not reject H0, There is no interaction between gender and age.

b) (B) Since p-value for gender is small, rates are affected by gender.

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: u1 = u2

Alternative hypothesis: At-least one of the u is not equal.

Formulate an analysis plan. For this analysis, the significance level is 0.05.

Analyze sample data.

F statistics is given by:-

F=\frac{MSB}{MSE}

F = 5.36

Fcritical = 4.41

The P-value = 0.033

Interpret results. Since the P-value (0.033) is less than the significance level (0.05), we have to reject the null hypothesis.

Conclusion:-

Reject H0, There is sufficient evidence for significant differences between the genders.

c) (A) Since p-value for age is small, rates are affected by age.

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: u1 = u2 = u3 = u4

Alternative hypothesis: At-least one of the u is not equal.

Formulate an analysis plan. For this analysis, the significance level is 0.05.

Analyze sample data.

F statistics is given by:-

F=\frac{MSB}{MSE}

F = 5.06

Fcritical = 3.56

The P-value = 0.018

Interpret results. Since the P-value (0.018) is less than the significance level (0.05), we have to reject the null hypothesis.

Conclusion:-

Reject H0, There is sufficient evidence for significant differences between the age groups.